OpenStudy (anonymous):

The cone pictured has a surface area of a0 square centimeters. (Use 3.14 for π.)

OpenStudy (anonymous):



OpenStudy (anonymous):

1134 ?

OpenStudy (anonymous):

I'm going to assume you do NOT want the calculus method shown for how the formula comes about and just give you the formula itself. Surface Area of Cone = \(\large\pi r s + \pi r^2\) Volume of Cone = \(\large\pi r^2 h\)